OpenStudy (anonymous):

What is the 10th term of an geometric sequence, 729, 243 81 27. express the answer in fraction

OpenStudy (anonymous):

the pattern here is division by three

OpenStudy (anonymous):

or to keep witht he definition of geometric sequences, multiplying by 1/3

OpenStudy (anonymous):

here r=1/3 and a=273 and n=is number of terms here is the formula an=ar^n